The following table gives the bushels of corn per acre, y resulting from the use of the various amounts of fertilizer in pounds per acre, x, produced on a farm in each of 10 years from 1971 to 1980.  

Year

n

y

x

1971

1

40

6

1972

2

44

10

1973

3

46

12

1974

4

48

14

1975

5

52

16

1976

6

58

18

1977

7

60

22

1978

8

68

24

1979

9

74

26

1980

10

80

32

A) Calculate the SST, SSE, and SSR of the model.

B) Calculate the estimated variances and the standard errors of and .

C) Calculate the R2 term of the estimated model and interpret the R2 term.
Answer #1
X Y (x-x̅)² (y-ȳ)² (x-x̅)(y-ȳ)
6 40 144 289 204
10 44 64 169 104
12 46 36 121 66
14 48 16 81 36
16 52 4 25 10
18 58 0 1 0
22 60 16 9 12
24 68 36 121 66
26 74 64 289 136
32 80 196 529 322
ΣX ΣY Σ(x-x̅)² Σ(y-ȳ)² Σ(x-x̅)(y-ȳ)
total sum 180 570 576 1634 956
mean 18 57 SSxx SSyy SSxy

sample size ,   n =   10      
here, x̅ =   18       ȳ =   57
              
SSxx =    Σ(x-x̅)² =    576      
SSxy=   Σ(x-x̅)(y-ȳ) =   956      
a)

SSE=   (Sx*Sy - S²xy)/Sx =    47.31
SSR=   S²xy/Sxx =   1586.69

SST=SSE+SSR=47.31+1586.69=1634.00

b)estimated variance=SSE/(n-2)=5.91319

estimate of std error=Se =    √(SSE/(n-2)) =    2.4317

c)

R² =    (Sxy)²/(Sx.Sy) =    0.9710
97.10% variation in observation of variable y, is explained by variable x.